# Learning to Estimate Critical Gait Parameters from Single-View RGB Videos with Transformer-Based Attention Network

### 01 Dec 2023 v1

@@ 1 Introduction
Recently, 2D markerless video sequences captured from single-camera videos have emerged as a convenient and economical option [11] by minimizing the needs for expensive equipment. In particular, deep neural networks (DNNs) become a powerful tool to automatically learn and extract 2D skeletal data [12] for human gait assessment [13]. However, estimating gait parameters directly from these skeletal data is prone to errors due to the lack of robustness and precision from pose tracking algorithms [14]. As a result, the expressiveness and resilience to noise of DNNs help them stand out as a more efficient and precise method for estimating general gait parameters [15]. In this study, we propose a novel spatio-temporal Transformer architecture for gait analysis using single-camera RGB videos, which could effectively learn intricate spatio-temporal dynamics in human gait, a challenge for most traditional methods. The first stage of this architecture aims to extract the keypoints of 2D human pose captured from a video using OpenPose [12]. The second stage uses a spatio-temporal Transformer network to predict gait parameters, including Gait Deviation Index, knee flexion angle at maximum extension, speed, and cadence. The main contributions of this work are summarized as follows:
  • We are the first to introduce a spatio-temporal Transformer network for quantitatively estimating gait parameters for gait analysis from single-view RGB videos. Our network design allows us to both extract spatial and temporal features effectively the anatomical keypoints from single-camera RGB videos.
  • The flexibility inherent in the proposed network help to achieve superior performance in predicting gait parameters compared to current state-of-the-art methods, meanwhile reducing number of parameters and obviating the need for manually crafted time series.
@@ 2 Related Works
Estimating critical gait parameters directly from single-camera RGB videos has been proposed in previous studies [15, 16, 17, 18]. For instance, Azhand et al. [16] extracted both 2D and 3D skeleton from a 2D input video and optimized the 3D pose based on patient’s heights to quantify gait parameters. Following the success of CNN in vision tasks, Kidzinski et al. [15] proposed to use OpenPose [12] for keypoint extraction and a one-dimensional convolutional neural network (1D-CNN) model to predict clinically relevant motion parameters from an ordinary video of a patient. Similarly, Lonini et al. [17] explored 1D-CNN to estimate gait parameters while utilizing DeepLabCut [19] for tracking anatomical keypoints of stroke patients. Researchers also employ Spatial-Temporal Graph Convolutional Network (STGCN) for gait parameter prediction [20]. For the graph of skeletal sequences, the connections of the joints are based on anatomical structure while every joint is connected to itself in the previous and the next frame. Transformer architecture is also proposed to compute interpretable kinematic parameters from estimated 3D pose, which are then used to estimate patients’ gait cycle and produce final gait parameters [18].
@@ 3.2 Network Architecture
We designed a Transformer-based network to estimate critical gait parameters directly from single-camera RGB videos. The network comprises of two main blocks: spatial attention and temporal attention blocks. This architecture is motivated by the success of the architecture proposed in [23] for three-dimensional pose estimation. The key idea behind self-attention is it enables the network to weigh the importance of every other element in a sequence when encoding each element, taking into account the relationships between them. In the spatial attention block, it determines the amount of focus of one particular joint has on every other joint in a single frame. This enables the model to learn the necessary spatial information between the joints.
@@
For each attention block, we use multi-head attention mechanism. In a multi-head attention block, multiple attention heads are calculated in parallel and independently. The output of this layer will be the concatenation of the outputs of the $H$ heads {IEEEeqnarray}rCl Attention(Q, K, V) &= Softmax(QK^T/D)V
@@ 4.1 Datasets and Experimental Settings
We validated our approach on a publicly available dataset collected at Gillette Children’s Specialty Healthcare between 1994 and 2015. Following preprocessing steps described in [15], the dataset consists of 2,212 videos of 1,138 patients, who had been diagnosed with cerebral palsy. The average age of the patients is 13 years old with standard deviation of 6.9 years old. The average height and weight of the patients are 141cm and 41kg with standard deviation of 19cm and 17kg, respectively. The raw videos were initially captured at 1280$\times$960 resolution and then downsampled to 640$\times$480 in order to better accommodate the capability of most low-end cameras. Each video is then divided into multiple 124-frame segments by window slicing with an offset of 31 frames. For experiments, the dataset is split into training, validation, and test set with the ratio of 8:1:1. It is divided by patients’ ID to ensure no patient is included in different set. The training set consists of 1,768 videos of 920 patients, meanwhile the validation set consists of 212 videos of 106 patients, and the test set consists of 232 videos of 112 patients.

### 01 Mar 2024 v2

@@ 1 Introduction
Recently, 2D markerless video sequences captured from single-camera videos have emerged as a convenient and economical option [11] by minimizing the needs for expensive equipment. In particular, pose estimation models become a powerful tool to automatically learn and extract 2D skeletal data [12] for human gait assessment [13]. However, estimating gait parameters directly from these skeletal data is prone to errors due to the lack of robustness and precision from pose tracking algorithms [14]. As a result, the expressiveness and resilience to noise of deep neural networks enable them to be a more efficient and precise method for estimating general gait parameters [15]. In this study, we propose a novel spatio-temporal Transformer architecture for gait analysis using single-camera RGB videos, which could effectively learn intricate spatio-temporal dynamics in human gait, a challenge for most traditional methods. The first stage of this architecture aims to extract the keypoints of 2D human pose captured from a video using OpenPose [12]. The second stage uses a spatio-temporal Transformer network to predict gait parameters, including Gait Deviation Index, knee flexion angle at maximum extension, speed, and cadence. The main contributions of this work are summarized as follows:
  • We are the first to introduce a spatio-temporal Transformer network for quantitatively estimating gait parameters for gait analysis from single-view RGB videos. Our network design allows us to extract both spatial and temporal features effectively from the anatomical keypoints taken from single-camera RGB videos.
  • The flexibility inherent in the proposed network helps to achieve superior performance in predicting gait parameters compared to current state-of-the-art methods, meanwhile reducing number of parameters and obviating the need for manually crafted time series.
@@ 2 Related Works
Estimating gait parameters directly from single-camera RGB videos has been proposed in previous studies [15, 16, 17, 18]. For instance, Azhand et al. [16] extracted both 2D and 3D skeleton from a 2D input video and optimized the 3D pose based on patient’s heights to quantify gait parameters. Following the success of CNN in vision tasks, Kidzinski et al. [15] proposed to use OpenPose [12] for keypoint extraction and a one-dimensional convolutional neural network (1D-CNN) model to predict clinically relevant motion parameters from an ordinary video of a patient. Similarly, Lonini et al. [17] explored 1D-CNN to estimate gait parameters while utilizing DeepLabCut [19] for tracking anatomical keypoints of stroke patients. Researchers also employ Spatial-Temporal Graph Convolutional Network (STGCN) for gait parameter prediction [20]. For the graph of skeletal sequences, the connections of the joints are based on anatomical structure while every joint is connected to itself in the previous and the next frame. Transformer architecture is also proposed to compute interpretable kinematic parameters from estimated 3D pose, which are then used to estimate patients’ gait cycle and produce final gait parameters [18].
@@ 3.2 Network Architecture
We designed a Transformer-based network to estimate critical gait parameters directly from single-camera RGB videos. The network comprises of two main blocks: spatial attention and temporal attention blocks. This architecture is motivated by the success of the architecture proposed in [23] for three-dimensional pose estimation. The key idea behind self-attention is it enables the network to weigh the importance of every other element in a sequence when encoding each element, taking into account the relationships between them. In the spatial attention block, it determines the amount of focus of one particular joint has on every other joint in a single frame. This allows the model to learn the necessary spatial information between the joints.
@@
For each attention block, we use multi-head attention (MHA) mechanism. In a multi-head attention block, multiple attention heads are calculated in parallel and independently. The output of this layer will be the concatenation of the outputs of the $H$ heads {IEEEeqnarray}rCl Attention(Q, K, V) &= Softmax(QK^T/D)V
@@ 4.1 Datasets and Experimental Settings
We validated our approach on a publicly available dataset collected at Gillette Children’s Specialty Healthcare between 1994 and 2015. Following the preprocessing steps described in [15], the dataset consists of 2,212 videos of 1,138 patients, who had been diagnosed with cerebral palsy. The average age of the patients is 13 years old with standard deviation of 6.9 years old. The average height and weight of the patients are 141cm and 41kg with standard deviation of 19cm and 17kg, respectively. The raw videos were initially captured at 1280$\times$960 resolution and then downsampled to 640$\times$480 in order to better accommodate the capability of most low-end cameras. Each video is then divided into multiple 124-frame segments by window slicing with an offset of 31 frames. For experiments, the dataset is split into training, validation, and test set with the ratio of 8:1:1. It is divided by patients’ ID to ensure no patient is included in different set. The training set consists of 1,768 videos of 920 patients, meanwhile the validation set consists of 212 videos of 106 patients, and the test set consists of 232 videos of 112 patients.